# RSTAR4D: Rotational streak artifact reduction in 4D CBCT using a separable 4D CNN11footnotemark: 1

### v2

@@ Abstract
Four-dimensional cone-beam computed tomography (4D CBCT) provides respiration-resolved images and can be used for image-guided radiation therapy. However, the ability to reveal respiratory motion comes at the cost of image artifacts. As raw projection data are sorted into multiple respiratory phases, the cone-beam projections become much sparser and the reconstructed 4D CBCT images will be covered by severe streak artifacts. Although several deep learning-based methods have been proposed to address this issue, most algorithms employ 2D network models as backbones, neglecting the intrinsic structural priors within 4D CBCT images. In this paper, we first explore the origin and appearance of streak artifacts in 4D CBCT images. We find that streak artifacts exhibit a unique rotational motion along with the patient’s respiration, distinguishable from diaphragm-driven respiratory motion in the spatiotemporal domain. Therefore, we propose a novel 4D neural network model, RSTAR4D-Net, designed to address Rotational STreak Artifact Reduction by integrating the spatial and temporal information within 4D CBCT images. Specifically, we overcome the computational and training difficulties of a 4D neural network. The specially designed model adopts an efficient implementation of 4D convolutions to reduce computational costs and thus can process the whole 4D image in one pass. Additionally, a Tetris training strategy pertinent to the separable 4D convolutions is proposed to effectively train the model using limited 4D training samples. Extensive experiments substantiate the effectiveness of our proposed method, and the RSTAR4D-Net shows superior performance compared to other methods. The source code and dynamic demos are available at https://github.com/ivy9092111111/RSTAR.
@@
MSC:
@@
journal: Medical Image Analysis
@@ 1 Introduction
Despite the impressive abilities of data-driven algorithms in image recovery, existing 2D methods may not fully exploit the intrinsic structural priors of 4D CBCT images. Specifically, the unique imaging protocol of 4D CBCT generates a distinct dynamic pattern of streak artifacts (will be discussed in Section 2). The temporal features could be valuable for the network model to differentiate artifacts from desired anatomical structures. Additionally, respiratory motion, mainly driven by the moving diaphragm, causes notable image deformation along the z-axis. 2D methods neglect the inter-phase and inter-plane correlations among 4D CBCT images and focus solely on in-plane image enhancement, which makes it difficult to identify artifacts and may introduce inconsistencies in the recovered image.
In this paper, we first investigate the origin and characteristics of streak artifacts in 4D CBCT images. We find that phase-specific streak artifacts manifest in a distinctive rotational pattern within a breathing cycle. Intuitively, we show that the motion pattern of the streak artifacts differs from that of respiratory motion. This dissimilarity suggests the possibility of decoupling them in the spatiotemporal space. Building upon this insight, we propose a 4D neural network model named RSTAR4D-Net for Rotational STreak Artifact Reduction in 4D CBCT. Unlike conventional approaches, the RSTAR4D-Net takes the whole 4D CBCT image as the input and can concurrently consider the spatial and temporal information. Although the idea is straightforward, we argue it is nontrivial to develop a 4D network model due to higher computational costs and the demand for large-scale 4D datasets. The main methodological innovation of this paper lies in the way we instantiate the 4D network model. Specifically, the RSTAR4D-Net implements separable 4D convolutions by dividing the convolutional operation into separate dimensions, making it efficient enough to process the whole 4D image in one pass. Additionally, the specially designed model allows us to adopt a novel Tetris training strategy to effectively train the 4D model with very limited 4D samples. Experimental results on both simulated and real clinical data confirm the superior performance of our method.
2 Method
@@ 2.1 Rotational streak artifacts (RSA) in 4D CBCT images
Fig. 1: 4D CBCT imaging with breathing signals. (a) The average image with blurry structure edges, (b) the breathing signal used to sort the projection data into different phases, (c) the phase sorting map, and (d) a phase-resolved image reconstructed from single-phase projection data, which better reveals the edge of the diaphragm but suffers severe streak artifacts.
@@
Fig. 2: Illustration of rotational streak artifacts (RSA). (a1)-(a5) The projection sampling patterns at five respiratory phases, (b1)-(b5) phase-resolved images with RSA.
@@
Fig. 3: Motion trajectories of densely sampled feature points projected to axial and coronal planes. (a) Respiration-induced motion in the axial plane, (b) RSA-induced motion in the axial plane, (c) Hybrid motion in the axial plane, (d) t-SNE plot of the three types of motion trajectory; (e) Respiration-induced motion in the coronal plane, (f) RSA-induced motion in the coronal plane, (g) Hybrid motion in the coronal plane, (h) The average optical flow component along different axes.
@@ 2.2 RSTAR4D-Net
Fig. 4: The design of the RSTAR4D-Net. (a) The RSTAR4D-Net takes a whole 4D CBCT image as the input. (b1-b2) 4D convolutional modules: (b1) Vanilla isotropic 4D convolutions, (b2) lightweight separable 4D convolutions (proposed).
@@ 2.2.1 Model Architecture
We leverage the well-known U-Net [20] as the backbone model. Initially, we replace the original 2D convolutional filters with 4D convolutional filters to accommodate the input data (see Fig. 4 (b1)). However, using isotropic 4D convolutions is computationally expensive. For example, a 3$\times$3$\times$3$\times$3 spatiotemporal convolutional filter requires nine times the parameters and computational costs compared to a 3$\times$3 filter. Instead, we decompose the 4D convolution into multiple lower-dimensional convolutions. Specifically, our implementation decomposes the 4D convolution into a 2D spatial convolution in the x-y plane, a 1D spatial convolution in the z dimension, and a 1D temporal convolution in the temporal dimension (see Fig. 4 (b2)). This approach increases the number of parameters and computational cost by only two-thirds while achieving a 4D convolution. Additionally, separable convolutions alleviate the training difficulty, which will be discussed in the next subsection. A similar idea has been explored to modify 3D convolutions in action recognition research for natural videos [23] and recently in large video generation models [2]. However, to our knowledge, this is the first attempt to develop a 4D neural network model for medical imaging research. With the separable convolution strategy, the RSTAR4D-Net can process a whole 4D CBCT image of size 256$\times$256$\times$90$\times$10 (height, width, slice, frame) in one pass at the inference stage with a single consumer-grade GPU (an Nvidia RTX 3090 with 24GB memory). Furthermore, instead of using standard zeros padding, we employ circular padding for the temporal convolution to better align with the quasi-periodic nature of respiratory motion. As a result, the proposed RSTAR4D-Net demonstrates a high capacity for recurrently extracting spatiotemporal information from the 4D CBCT image.
@@ 2.2.2 Tetris training strategy
Fig. 5: The Tetris training strategy. (a) The RSTAR4D-Net processes an entire 4D CBCT image at the inference stage. (b1) 2D+T image series used in Tetris training stage I, (b2) 4D image blocks of various sizes used in Tetris training stage II.
In Tetris stage I, we train the RSTAR4D-Net using 2D+T data. Specifically, we reformat the 4D CBCT image slice-by-slice (in axial view) to significantly increase the number of samples in the dataset. The separable convolutions in the RSTAR4D-Net include three relatively independent data flows to encode the image features along different dimensions. Thus, we can simply freeze and skip the 1D convolution module along the z-axis in Tetris stage I while maintaining the overall architecture of the model. As RSA primarily affects the axial plane and exhibits a unique motion pattern along the temporal dimension (as analyzed in Section 2.1), Tetris stage I enables the model to recognize and reduce the RSA effectively. Additionally, training on 2D+T data largely enhances data throughput and accelerates convergence.
In Tetris stage II, we finetune the RSTAR4D-Net using 4D image blocks of various sizes. The 4D CBCT images used in this study have a size of 256$\times$256$\times$90$\times$10 (height, width, slice, frame). In our implementation, we crop these 4D images into blocks of three different sizes: 256$\times$256$\times$16$\times$10, 144$\times$144$\times$32$\times$10, and 128$\times$128$\times$64$\times$10. During this stage, we unfreeze the 1D z-axis convolution module and optimize the entire 4D model on 4D samples. As the diaphragm-driven respiratory motion leads to image deformation mainly along the z-axis (as analyzed in Section 2.1), Tetris stage II allows the model to accommodate respiratory motion and enhance volumetric consistency. We avoid training directly on the entire 4D CBCT image (size of 256$\times$256$\times$90$\times$10) due to high memory requirements. Additionally, the cropping step serves as data augmentation, helping to mitigate the overfitting problem (the number of training samples used in this study will be shown in Table 1).
@@ 3.1.1 Simulated training and validation datasets
3.1.2 Simulated test dataset
3.1.3 Real clinical test dataset
3.2 Comparison Methods
3.3 Implementation Details
4 Results
4.1 Experiments on the simulated dataset
@@
Fig. 6: Axial reconstruction results of the simulated dataset arranged as follows: (a) Reference, (b) Gated FDK, (c) PICCS, (d) MSD-GAN, (e) CycN-Net, (f) PriorNet, (g) RSTAR4D-Net. The first and the third rows present the reconstructed images with zoom-in tumor regions, and the second and the fourth rows present the corresponding difference images in absolute value. The display window is [-1000, 500] HU for reconstructed images, [-450, 150] HU for zoom-in tumor regions, and [0, 300] HU for absolute difference images.
Fig. 7: Coronal reconstruction results of the simulated dataset arranged as follows: (a) Reference, (b) Gated FDK, (c) PICCS, (d) MSD-GAN, (e) CycN-Net, (f) PriorNet, (g) RSTAR4D-Net. The first and the third rows present the reconstructed images with zoom-in tumor regions, and the second and the fourth rows present the corresponding difference images in absolute value. The display window is [-1000, 500] HU for reconstructed images, [-450, 150] HU for zoom-in tumor regions, and [0, 300] HU for absolute difference images.
@@
Fig. 8: Reconstruction results of the real clinical dataset (cases 1-3) in coronal views arranged as follows: (a) Gated FDK, (b) PICCS, (c) MSD-GAN, (d) CycN-Net, (e) PriorNet, (f) RSTAR4D-Net. The display window is [-1000, 500] HU for reconstructed images. The red boxes indicate the tumors. The horizontal yellow lines indicate the diaphragms. The yellow arrows denote the volumetric inconsistencies along the z-axis.
@@ 4.2 Experiments on the real clinical dataset
Fig. 9: Comparison between the real and recovered projection data (clinical case 4). (a) Real Projection data. (b-e) SSIM maps of different methods: (b) MSD-GAN, (c) CycN-Net, (d) PriorNet, (e) RSTAR4D-Net.
In this section, we present the results of applying our RSTAR4D-Net to real clinical datasets, showcasing its performance in a practical setting. In Fig. 8, we present the reconstructed results of three real clinical CBCT scans (cases 1, 2, and 3). Although ground truth images are unavailable, we evaluate the performance of each method by analyzing the positions of the tumor and diaphragm. The gated FDK algorithm reconstructs phase-resolved images using only the projection data collected at the specific phase, resulting in severe streak artifacts due to insufficient sampling. However, the degraded images still help locate the position of the moving tumor and diaphragm (please see Fig.8 (a1)-(a6)), as inter-phase interference is omitted. Thus, we take the gated FDK images as a reference to determine if the methods correctly removed streak artifacts while preserving anatomical structures. For each sub-figure, the tumor is denoted with a red box, and the relative position of each red box remains the same. Horizontal yellow lines mark the diaphragm’s position. While all deep learning-based methods reduce streak artifacts on real clinical 4D CBCT images, Fig. 8 reveals that anatomical structures may be inaccurately estimated by the comparison methods. Specifically, volumetric inconsistency is observed in the diaphragm regions (noted by the yellow arrows). Additionally, a lower diaphragm’s position is observed at the end-exhale phase (please see Fig. 8 (d3), (e3), (c5)), while a higher diaphragm’s position is observed at the end-inhale phase (please see Fig. 8 (e2), (d6), (e6)). These results suggest that the comparison methods may overly rely on the average image, failing to estimate the exact amplitude of respiratory motion. In contrast, the proposed RSTAR4D-Net successfully recovers the positions of the moving organs.
@@
5 Conclusion and Discussion
In this paper, we proposed a novel deep learning-based method for reducing rotation streak artifacts (RSA) in 4D CBCT. Our approach integrates domain knowledge into a data-driven framework. Initially, we investigated the origin and appearance of streak artifacts in 4D CBCT images. Thereafter, we identified the unique motion pattern of RSA, motivating us to decouple the artifacts from the anatomical structures in the spatiotemporal domain. Subsequently, we introduced the RSTAR4D-Net to recover the 4D CBCT images. The RSTAR4D-Net leverages separable 4D convolution and a temporally circular padding strategy to adapt to the physiological nature of respiratory motion. Extensive experiments validate the effectiveness of the proposed method. The proposed RSTAR4D-Net is a lightweight and flexible model, we posit that incorporating advanced modules (e.g., temporal transformer [5], spatiotemporal implicit neural representation [8]) may further enhance the performance. Furthermore, we proposed the Tetris training strategy to effectively train a 4D neural network model with only a few 4D training samples. We believe the proposed method could be generalized to various 4D imaging scenarios, such as dynamic MR, PET, and SPECT imaging.

### v4

# RSTAR4D: Rotational Streak Artifact Reduction in 4D CBCT Using Separable 4D Convolutions

@@ Abstract
Four-dimensional cone-beam computed tomography (4D CBCT) provides respiration-resolved images and can be used for image-guided radiation therapy. However, the ability to reveal respiratory motion comes at the cost of image artifacts. As raw projection data are sorted into multiple respiratory phases, the cone-beam projections become much sparser and the reconstructed 4D CBCT images are covered by severe streak artifacts. Although several deep learning-based methods have been proposed to address this issue, most algorithms formulate it as a 2D image enhancement task, neglecting the dynamic nature of 4D CBCT images. In this paper, we first identify the origin and appearance of streak artifacts in 4D CBCT images. We find that streak artifacts exhibit a unique “rotational motion” along with the patient’s respiration, distinguishable from diaphragm-driven respiratory motion in 4D space. Therefore, we propose a novel 4D neural network model, RSTAR4D-Net, designed to address Rotational STreak Artifact Reduction by exploring the dynamic prior of 4D CBCT images. Specifically, we overcome the computational and training difficulties of a 4D neural network. The specially designed model decomposes the 4D convolutions into multiple lower-dimensional operations and thus can efficiently excavate the spatiotemporal information from a whole 4D image. Additionally, a Tetris training strategy pertinent to the separable 4D convolutions is proposed to effectively train the model using limited 4D training samples. Extensive experiments substantiate the effectiveness of our proposed method, and the RSTAR4D-Net shows superior performance compared to other methods. The source code and dynamic demos are available at https://github.com/ivy9092111111/RSTAR.
@@
Computed tomography , Motion artifact reduction , Deep learning , 4D imaging
journal: Computerized Medical Imaging and Graphics
@@
[label1]organization=School of Biomedical Engineering, Shanghai Jiao Tong University,city=Shanghai, country=China
[…]
[label2]organization=Department of Radiation Oncology, Shanghai Chest Hospital, Shanghai Jiao Tong University School of Medicine,city=Shanghai, country=China
[…]
[label3]organization=Department of Respiratory Endoscopy, Department of Respiratory and Critical Care Medicine, Shanghai Chest Hospital, Shanghai Jiao Tong University School of Medicine,city=Shanghai, country=China
[…]
[label4]organization=Jiangsu First-Imaging Tech.,city=Shanghai, country=China
[…]
[label5]organization=Research Center of Augmented Intelligence, Zhejiang Lab,city=Hangzhou, country=China
[…]
[label6]organization=School of Computer Science and Engineering, Southeast University,city=Nanjing, country=China
@@ 1 Introduction
Despite the impressive abilities of data-driven algorithms in image recovery, existing 2D methods may not fully exploit the intrinsic structural priors of 4D CBCT images. Specifically, the unique imaging protocol of 4D CBCT generates a distinct dynamic pattern of streak artifacts (will be discussed in 2). The temporal features could be valuable for the network model to differentiate artifacts from desired anatomical structures. Additionally, respiratory motion, mainly driven by the moving diaphragm, causes notable image deformation along the z-axis. 2D methods neglect the inter-phase and inter-plane correlations among 4D CBCT images and focus solely on in-plane image enhancement, which makes it difficult to identify artifacts and may introduce inconsistencies in the recovered image.
In this paper, we first investigate the origin and characteristics of streak artifacts in 4D CBCT images. We find that phase-specific streak artifacts manifest in a distinctive rotational pattern within a breathing cycle. Intuitively, we show that the motion pattern of the streak artifacts differs from that of respiratory motion. This dissimilarity suggests the possibility of decoupling them in 4d space. Building upon this insight, we propose a 4D neural network model named RSTAR4D-Net for Rotational STreak Artifact Reduction in 4D CBCT. Unlike conventional approaches, the RSTAR4D-Net takes the whole 4D CBCT image as the input and can concurrently consider the spatial and temporal information. Although the idea is straightforward, we argue it is nontrivial to develop a 4D network model due to higher computational costs and the demand for large-scale 4D datasets. The main methodological innovation of this paper lies in the way we instantiate the 4D network model. Specifically, the RSTAR4D-Net implements separable 4D convolutions by dividing the convolutional operation into separate dimensions, making it efficient enough to process the whole 4D image in one pass. Additionally, the specially designed model allows us to adopt a novel Tetris training strategy to effectively train the 4D model with very limited 4D samples. Experimental results on both simulated and real clinical data confirm the superior performance of our method.
Figure 1: 4D CBCT imaging with breathing signals. (a) The average image with blurry structure edges, (b) the breathing signal used to sort the projection data into different phases, (c) the phase sorting map, and (d) a phase-resolved image reconstructed from single-phase projection data, which reveals the edge of the diaphragm but suffers severe streak artifacts.
2 Material and methods
@@ 2.1 Rotational Streak Artifacts (RSA) in 4D CBCT Images
Figure 2: Illustration of rotational streak artifacts (RSA). (a1)-(a5) The projection sampling patterns at five respiratory phases, (b1)-(b5) phase-resolved images with RSA.
@@
Figure 3: Motion trajectories of densely sampled feature points projected to axial and coronal planes. (a) Respiration-induced motion in the axial plane, (b) RSA-induced motion in the axial plane, (c) Hybrid motion in the axial plane, (d) t-SNE plot of the three types of motion trajectory; (e) Respiration-induced motion in the coronal plane, (f) RSA-induced motion in the coronal plane, (g) Hybrid motion in the coronal plane, (h) The average optical flow component along different axes.
Figure 4: The design of the RSTAR4D-Net. (a) The RSTAR4D-Net takes a whole 4D CBCT image as the input. (b1-b2) 4D convolutional modules: (b1) Vanilla isotropic 4D convolutions, (b2) lightweight separable 4D convolutions (proposed).
@@ 2.2.1 Model Architecture
We leverage the well-known U-Net [25] as the backbone model. Initially, we replace the original 2D convolutional filters with 4D convolutional filters to accommodate the input data (see Fig. 4 (b1)). However, using isotropic 4D convolutions is computationally expensive. For example, a 3$\times$3$\times$3$\times$3 spatiotemporal convolutional filter requires nine times the parameters and computational costs compared to a 3$\times$3 filter. Instead, we decompose the 4D convolution into multiple lower-dimensional convolutions. Specifically, our implementation decomposes the 4D convolution into a 2D spatial convolution in the x-y plane, a 1D spatial convolution in the z dimension, and a 1D temporal convolution in the temporal dimension (see Fig. 4 (b2)). This approach increases the number of parameters and computational cost by only two-thirds while achieving a 4D convolution. A similar idea has been explored to modify 3D convolutions in action recognition research for natural videos [26] and recently in large video generation models [27]. However, to our knowledge, this is the first attempt to develop a 4D neural network model for 4D CBCT research. Additionally, separable convolutions alleviate the training difficulty, which will be discussed in the next subsection. With the separable convolution strategy, the RSTAR4D-Net can process a whole 4D CBCT image of size 256$\times$256$\times$90$\times$10 (height, width, slice, frame) in one pass at the inference stage with a single consumer-grade GPU (an Nvidia RTX 3090 with 24GB memory). Furthermore, instead of using standard zeros padding, we employ circular padding for the temporal convolution to better align with the quasi-periodic nature of respiratory motion. As a result, the proposed RSTAR4D-Net demonstrates a high capacity for recurrently extracting spatiotemporal information from the 4D CBCT image.
@@ 2.2.2 Tetris training strategy
Figure 5: The Tetris training strategy. (a) The RSTAR4D-Net processes an entire 4D CBCT image at the inference stage. (b1) 2D+T images used in Tetris training stage I, (b2) 4D image blocks of various sizes used in Tetris training stage II.
In Tetris stage I, we train the RSTAR4D-Net using 2D+T data. Specifically, we reformat the 4D CBCT image slice-by-slice (in axial view) to significantly increase the number of samples in the dataset. The separable convolutions in the RSTAR4D-Net include three relatively independent data flows to encode the image features along different dimensions. Thus, we can simply freeze and skip the 1D convolution module along the z-axis in Tetris stage I while maintaining the overall architecture of the model. As RSA primarily affects the axial plane and exhibits a unique motion pattern along the temporal dimension (as analyzed in 2.1), Tetris stage I enables the model to recognize and reduce the RSA effectively. Additionally, training on 2D+T data largely enhances data throughput and accelerates convergence.
In Tetris stage II, we finetune the RSTAR4D-Net using 4D image blocks of various sizes. The 4D CBCT images used in this study have a size of 256$\times$256$\times$90$\times$10 (height, width, slice, frame). In our implementation, we empirically crop these 4D images into blocks of three different sizes: 256$\times$256$\times$16$\times$10, 144$\times$144$\times$32$\times$10, and 128$\times$128$\times$64$\times$10. During this stage, we unfreeze the 1D z-axis convolution module and optimize the entire 4D model on 4D samples. As the diaphragm-driven respiratory motion leads to image deformation mainly along the z-axis (as analyzed in 2.1), Tetris stage II allows the model to accommodate respiratory motion and enhance volumetric consistency. We avoid training directly on the entire 4D CBCT image (size of 256$\times$256$\times$90$\times$10) due to high memory requirements. Additionally, the cropping step serves as data augmentation, helping to mitigate the overfitting problem.
@@ 2.3.1 Simulated training and validation datasets
2.3.2 Simulated test dataset
2.3.3 Real clinical test dataset
2.3.4 Comparison Methods
2.3.5 Implementation Details
3 Results
3.1 Experiments on the Simulated Dataset
@@
Figure 6: Axial reconstruction results of the simulated dataset arranged as follows: (a) Reference, (b) Gated FDK, (c) PICCS, (d) MSD-GAN, (e) CycN-Net, (f) PriorNet, (g) RSTAR4D-Net. The first and the third rows present the reconstructed images with zoom-in tumor regions, and the second and the fourth rows present the corresponding difference images in absolute value. The display window is [-1000, 500] HU for reconstructed images, [-450, 150] HU for zoom-in tumor regions, and [0, 300] HU for absolute difference images.
Figure 7: Coronal reconstruction results of the simulated dataset arranged as follows: (a) Reference, (b) Gated FDK, (c) PICCS, (d) MSD-GAN, (e) CycN-Net, (f) PriorNet, (g) RSTAR4D-Net. The first and the third rows present the reconstructed images with zoom-in tumor regions, and the second and the fourth rows present the corresponding difference images in absolute value. The display window is [-1000, 500] HU for reconstructed images, [-450, 150] HU for zoom-in tumor regions, and [0, 300] HU for absolute difference images.
@@ 3.2 Experiments on the Real Clinical Dataset
Figure 8: Reconstruction results of the real clinical dataset (cases 1-3) in coronal views arranged as follows: (a) Gated FDK, (b) PICCS, (c) MSD-GAN, (d) CycN-Net, (e) PriorNet, (f) RSTAR4D-Net. The display window is [-1000, 500] HU for reconstructed images. The red boxes indicate the tumors. The horizontal yellow lines indicate the diaphragms. The yellow arrows denote the volumetric inconsistencies along the z-axis.
In this section, we present the results of applying our RSTAR4D-Net to real clinical datasets, showcasing its performance in a practical setting. In Fig. 8, we present the reconstructed results of three real clinical CBCT scans (cases 1, 2, and 3). Although ground truth images are unavailable, we evaluate the performance of each method by analyzing the positions of the tumor and diaphragm [31]. The gated FDK algorithm reconstructs phase-resolved images using only the projection data collected at the specific phase, resulting in severe streak artifacts due to insufficient sampling. However, the degraded images still help locate the position of the moving tumor and diaphragm (please see Fig.8 (a1)-(a6)), as inter-phase interference is omitted. Thus, we take the gated FDK images as a reference to determine if the methods correctly removed streak artifacts while preserving anatomical structures. For each sub-figure, the tumor is denoted with a red box, and the relative position of each red box remains the same. Horizontal yellow lines mark the diaphragm’s position. While all deep learning-based methods reduce streak artifacts on real clinical 4D CBCT images, Fig. 8 reveals that anatomical structures may be inaccurately estimated by the comparison methods. Specifically, volumetric inconsistency is observed in the diaphragm regions (noted by the yellow arrows). Additionally, a lower diaphragm’s position is observed at the end-exhale phase (please see Fig. 8 (d3), (e3), (c5)), while a higher diaphragm’s position is observed at the end-inhale phase (please see Fig. 8 (e2), (d6), (e6)). These results suggest that the comparison methods may overly rely on the average image, failing to estimate the exact amplitude of respiratory motion. In contrast, the proposed RSTAR4D-Net successfully recovers the positions of the moving organs.
Figure 9: Comparison between the real and recovered projection data (clinical case 4). (a) Real Projection data. (b-e) SSIM maps of different methods: (b) MSD-GAN, (c) CycN-Net, (d) PriorNet, (e) RSTAR4D-Net.
@@
4 Discussion and Conclusion
In this paper, we proposed a novel deep learning-based method for reducing rotation streak artifacts (RSA) in 4D CBCT. Our approach integrates domain knowledge into a data-driven framework. Initially, we investigated the origin and appearance of streak artifacts in 4D CBCT images. Thereafter, we identified the unique motion pattern of RSA, motivating us to decouple the artifacts from the anatomical structures in the spatiotemporal domain. Subsequently, we introduced the RSTAR4D-Net to recover the 4D CBCT images. The RSTAR4D-Net leverages separable 4D convolution and a temporally circular padding strategy to adapt to the physiological nature of respiratory motion. Extensive experiments validate the effectiveness of the proposed method. The proposed RSTAR4D-Net is a lightweight and flexible model, we posit that incorporating advanced modules (e.g., temporal transformer [32], spatiotemporal implicit neural representation [33]) may further enhance the performance. Furthermore, we proposed the Tetris training strategy to effectively train a 4D neural network model with only a few 4D training samples. We believe the proposed method could be generalized to various 4D imaging techniques, such as dynamic MR, PET, and SPECT imaging.
With our method, during a 1-minute scan, clear 4D CBCT images can be obtained without the need for breath holding. The 4D CBCT images enable physicians to better understand the respiratory motion of the lesions and facilitate precise lesion localization during radiation therapy. The proposed method may bring along many new application scenarios. For example, the image-guided percutaneous lung biopsy is performed using CBCT. To acquire clear CBCT image, general anesthesia is needed for better respiratory control. With the proposed method, it may be possible for free-breathing CBCT scan and the procedures can be performed under moderate sedation with local anesthesia.